meganmoney543:

+ 3 = 27 solve for x.

TETSXPREME:

Are we supposed to simplify?

meganmoney543:

yes

TETSXPREME:

Ok so we basically divide 27 to 3

xxDeppressionxx:

Excuse me are there any more numbers to this equation

meganmoney543:

yes hol up

TETSXPREME:

see what number times 3 equals 27

TyHawkins:

use.. www.mathpapa.com

meganmoney543:

3x/4+3=27

TETSXPREME:

i thought you said we was supposed to simplify

meganmoney543:

i dunno i asked yall for help

xxDeppressionxx:

You did say we had to simplify right?

meganmoney543:

yee

TETSXPREME:

ok so then we just figure out what 3 Xs what makes 27

TETSXPREME:

or basically divide 27 by 3

meganmoney543:

yea ig

AZ:

This is not how you solve an equation

AZ:

We're solving for `x` We have to try to get `x` all by itself \(\dfrac{3x}{4} + 3 = 27\) First, we want to move 3 to the other side. What do you get when you subtract 3 on both sides?

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @TETSXPREME i think the answer is 9 idrk \(\color{#0cbb34}{\text{End of Quote}}\) This is not the answer either. @meganmoney543 Let's solve it step by step. What is 27-3 = ?

meganmoney543:

idk

meganmoney543:

AZ im to slow for the step by step bs

AZ:

It's the only way to learn, I'm afraid

AZ:

Can you count down on your fingers? 27 - 3 = ?

meganmoney543:

ight ima delete the question

XxChampion23Xx:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @meganmoney543 ight ima delete the question \(\color{#0cbb34}{\text{End of Quote}}\) close*

meganmoney543:

yes r u the person that corrects the words

xxDeppressionxx:

Subtract 3 from both sides of the equation 3π‘₯4+3=27 3π‘₯/4+3βˆ’3=27βˆ’3 Subtract the numbers 3π‘₯/4=24 Multiply all terms by the same value to eliminate fraction denominators 3π‘₯4=244β‹…3π‘₯/4=4β‹…24 Simplify Cancel multiplied terms that are in the denominator Multiply the numbers 3π‘₯=96 Divide both sides of the equation by the same term 3π‘₯=96 3π‘₯/3=96/3 Cancel terms that are in both the numerator and denominator Divide the numbers π‘₯=32 Solution π‘₯=32

meganmoney543:

kk thxx😁

xxDeppressionxx:

Np

meganmoney543:

can we be friends?

supie:

\(\LARGE\frac{3x}{4}+3=27 \longrightarrow \frac{3}{4}x+3=27\) The first step is to get `+3` then change the plus `(+)` to a minus `(-)` so that we now have `-3` Now that we have `-3` we have to subtract 3 from both sides of the equation Which should give us \(\LARGE\frac{3}{4}x=24\) Because \(\frac{3}{4}x+3-3=27-3\) (27-3=24) So we get \(\LARGE\frac{3}{4}x=24\) Then multiply both sides of the equation by \(\frac{4}{3}\) \(\LARGE\frac{4}{3} \times \frac{3}{4}x = \frac{4}{3} * 24\) \(\frac{4}{3} \times 24 =? \) \(x=?\)

meganmoney543:

what?

xxDeppressionxx:

Supie i already handled it

supie:

Oh o: I didn't see

meganmoney543:

lol

AZ:

It does feel nice to know that out of the 13 people who were viewing this post, there are at least two others who know how to solve an equation. (Although, @xxDeppressionxx 's response seems to be copy pasted from a calculator.)

Laylalyssa:

imagine copy pasting from a calculator

TETSXPREME:

sick

xxDeppressionxx:

How does everyone seem to find something to argue about?

Laylalyssa:

How does someone copy paste from a calculator and expect nobody to know?

KjSaif:

Ah, looks like I joined in on this a bit late. @meganmoney543 don't forget to close the question. Also, tag me if you need anymore help lol.

xxDeppressionxx:

How do you think arguing will help any of this?

TETSXPREME:

crac kheadz o;

xxDeppressionxx:

Prove that i did that.

Laylalyssa:

It's obvious. Are you denying that you copied from a calculator?

Laylalyssa:

For example this `3π‘₯4=244β‹…3π‘₯/4=4β‹…24` what does that even mean?

AZ:

Let's take this to the chats

smackz:

take 3 divide it by 27 whats your answer u get if u dont know its 9

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @smackz take 3 divide it by 27 whats your answer u get if u dont know its 9 \(\color{#0cbb34}{\text{End of Quote}}\) \(\color{#0cbb34}{\text{Originally Posted by}}\) @smackz take 3 divide it by 27 whats your answer u get if u dont know its 9 \(\color{#0cbb34}{\text{End of Quote}}\) It has already been established that this is incorrect.

smackz:

how? @AZ

AZ:

If you read through the entire post, you will find how to actually solve it :)

smackz:

erm i jus scrolled al lthe way down.

smackz:

ok lemem look for it roach ;o

xxDeppressionxx:

yes you might need to re-read the post to see that it is wrong.

xxDeppressionxx:

Please do not by-pass the filter.

AZ:

... because the answer doesn't come out to 24 as shown above?